OpenStudy (anonymous):

Which is a solution to the equation y = 2x – 5? A. (2, 1) B. (1, –3) C. (3, 11) D. (5, –5)

OpenStudy (anonymous):

r there to equations and u posted one?

OpenStudy (cgreenwade2000):

Plug in the value for x into the equation and see if you get the corresponding y value when you solve.

OpenStudy (anonymous):

Substitute the values given into the equation to see which one gives an equality. For example: y=2x-5 (2,1) 1=2(2)-5 1 = 4-5 1 doesn't equal -1 so try the b answer (1,-3) -3=2(1)-5 solve and see if both sides equal each other

OpenStudy (anonymous):

Plug in the answers in the equation: y = 2x – 5 (2, 1) 1 = 2(2) - 5 1 = 4 - 5 1 = -1 (1, -3) -3 = 2(1) - 5 -3 = 2 - 5 -3 = -3 (3, 11) 11 = 2(3) - 5 11 = 6 - 5 11 = 1 (5, -5) -5 = 2(5) - 5 -5 = 10 - 5 -5 = 5 So which is the correct answer?

OpenStudy (anonymous):

work it out

OpenStudy (anonymous):

@botdfgirl14